# Holonomy of the Obata connection on 2-step hypercomplex nilmanifolds

### v1

@@ Abstract
We study the holonomy of the Obata connection on 2-step hypercomplex nilmanifolds. By explicitly computing the curvature tensor, we determine the conditions under which the Obata connection is flat, showing that this depends on the nilpotency step of each complex structure. In particular, we show that the holonomy algebra of the Obata connection is always an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$: to establish this, we prove the $\mathbb{H}$-solvable conjecture for 2-step hypercomplex nilmanifolds. Furthermore, we provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
@@ 1 Introduction
Nilmanifolds, that is, compact quotients $\G\Gamma\backslash$ G of connected and simply connected nilpotent Lie groups $G$ by discrete subgroups $\Gamma$, have played a fundamental role in differential geometry and topology since the seminal works of Mal’cev and Mostow. They provide a natural class of spaces where algebraic and geometric properties interact tightly, offering a fair balance between algebraic manageability and rich geometric behavior. In low dimensions, the latter allows classification results for various geometric structures (e.g., [Sal, FPS, Ug]), especially when these are left-invariant.
@@
In the hypercomplex setting, many of these nice behaviors fail to carry over in a unified way. Each complex structure $J_{\alpha}$ ($\alpha=1,2,3$) may satisfy suitable compatibility conditions individually, but it is not clear whether these structures collectively induce analogous global properties. For instance, it is natural to ask whether the subspace
@@
A central geometric object in this setting is the Obata connection: the unique torsion-free connection preserving all three complex structures. Its holonomy group $\operatorname{Hol}(\nabla)$ lies in $\operatorname{GL}(n,\mathbb{H})$ and plays a crucial role in understanding the geometry of hypercomplex manifolds [IP, SV]. In general, explicit descriptions of $\operatorname{Hol}(\nabla)$ are rare (see, for instance, [Sol, AGT, BFGV]). For nilmanifolds, however, it is known that the holonomy is contained in the commutator subgroup $\operatorname{SL}(n,\mathbb{H})$ [BDV, Corollary 3.3].
@@
This naturally leads to the question: Are all 2-step nilpotent hypercomplex Lie algebras Obata-flat? We answer this in the negative, and clarify the role of the nilpotency step of each complex structure (see Sections 2 and 3 for details). In particular, we show that if the Obata connection is not flat, then each $J_{\alpha}$ must be 3-step nilpotent. No such examples had been previously known.
@@
In Section 6, we prove that the holonomy algebra of the Obata connection is always abelian for $2$-step nilpotent hypercomplex Lie groups, and we describe it explicitly (Theorem 6.1). This reveals a surprising rigidity: even when the Obata connection is non-flat, the holonomy remains severely restricted. Our approach relies on proving the Gorginyan conjecture (Conjecture 5.3) in the 2-step case (Theorem 5.4). We further show that the Obata holonomy group of a $2$-step hypercomplex nilmanifold is trivial if and only if the nilmanifold is a torus.
In the final part of the paper, we extend our analysis to the $3$-step nilpotent case, showing that, under suitable conditions, the holonomy of the Obata connection remains abelian.
Acknowledgments. A. Andrada and M. L. Barberis were partially supported by CONICET, SECyT-UNC and ANPCyT (Argentina). B. Brienza was partially supported by Project PRIN 2022 “Real and complex manifolds: Geometry and Holomorphic Dynamics”, by GNSAGA (Indam) and by the INdAM - GNSAGA Project CUP E53C24001950001. The third author is grateful to A. Fino, G. Gentili, G. Gorginyan, and M. Verbitsky for many helpful discussions, and to the FAMAF for the warm hospitality during the preparation of this work.
@@ 2 Preliminaries
It is known that $(M,\{J_{\alpha}\})$ is quaternionic, (i.e., locally modelled over $\mathbb{H}^{n}$) if and only if the Obata connection $\nabla$ is flat [Ob, Theorem 11.2]. This highlights a fundamental difference with the complex case. For instance, the $K3$ surfaces and the homogeneous spaces $\operatorname{SU}(3)$ and $\operatorname{SO}(6)/\operatorname{SU}(2)$ provide examples of hypercomplex but not quaternionic manifolds [Boy, Joy].
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A hypercomplex structure $\{J_{\alpha}\}$ on $G$ is said to be left invariant if left translations by elements of $G$ are hyperholomorphic, i.e., holomorphic with respect to each $J_{\alpha}$. In this case, the structure is completely determined by its value at the identity, which corresponds to a hypercomplex structure on $\mathfrak{g}$.
If $\Gamma$ is a lattice in $G$, any left invariant hypercomplex structure on $G$ descends to a hypercomplex structure on the compact quotient $\G\Gamma\backslash$ G, which is then called invariant. In this setting, the natural projection $\GG\to\Gamma\backslash G$ is hyperholomorphic.
@@ A Lie algebra $\mathfrak{g}$ is said to be nilpotent if its lower central series
It is well known that the existence of a lattice for a connected and simply connected nilpotent Lie group $N$ is equivalent to the existence of a rational basis of its Lie algebra $\mathfrak{g}$ [Mal].
@@
The ascending series $\{\mathfrak{a}_{k}\ ;\ k\geq 0\}$ (compatible with $J$) of $\mathfrak{g}$ is defined inductively by
$\mathfrak{a}_{0}=0,\ \mathfrak{a}_{k}=\{X\in\mathfrak{g}\ |\ [X,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\ \ \text{and}\ \ [JX,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\},\ k\geq 1.$ (2.4)
@@ 3 Computation of the Obata curvature
    where the second equality follows from the integrability of $J_{\alpha}$ on the first term and the integrability of $J_{\gamma}$ on the second term. ∎
As a consequence of Proposition 3.1 we obtain the following result. This is a generalization of [B1, Proposition 5.1], where it was shown that if $\mathfrak{z}$ is $J_{\alpha}$ invariant for all $\alpha$ then $\{J_{\alpha}\}$ is Obata flat.
@@
  • Proof.
    If $J_{\beta}$ is $2$-step nilpotent, $J_{\beta}\mathfrak{g}^{1}\subset\mathfrak{z}$ (Lemma 2.4). The result then follows from Proposition 3.1, since each term contains an element of the form $\cdot,J_{\beta}\mathfrak{g}^{1}]$. ∎
@@ 4 Construction of non-flat $2$-step hypercomplex nilmanifolds
To the best of our knowledge, in all known examples of $2$-step nilpotent hypercomplex Lie algebras $(\mathfrak{g},\{J_{\alpha}\})$, each complex structure $J_{\alpha}$ is $2$-step [B2, DF1, DF2, BD, LW, FG]. In particular, every known example of a $2$-step nilpotent Lie group equipped with a left-invariant hypercomplex structure is Obata flat (see Corollary 3.2).
According to [DF2, Section 3], in dimension $8$ any nilpotent Lie algebra admitting a hypercomplex structure is $2$-step nilpotent, and, moreover, each complex structure $J_{\alpha}$ is $2$-step. This naturally raises the question of whether this behavior is general among $2$-step nilpotent hypercomplex Lie algebras, or whether there exist examples in which one, two, or all of the complex structures are $3$-step.
Pushing this further: if such examples where each $J_{\alpha}$ is $3$-step exist, are the corresponding $2$-step nilpotent hypercomplex Lie groups still Obata flat? Looking at Formula (3.1), this is not clear.
In this section, we address these questions.
@@
Let $\mathfrak{n}$ be a $2$-step nilpotent Lie algebra. We consider $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{r}$.
Let $\mu\in\bigwedge^{2}\mathfrak{n}^{*}\otimes\mathbb{R}^{r}$ be a $2$-form on $\mathfrak{n}$ with values in $\mathbb{R}^{r}$ such that $\mathfrak{n}^{1}\subset\ker(\mu)$. We use $\mu$ to define a bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ as follows:
@@
Then, $(\mathfrak{g},[\cdot,\cdot]^{\prime})$ is a $2$-step nilpotent Lie algebra.
Let us now assume that $\mathfrak{n}$ is a $2$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$, and let $(\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k},[\cdot,\cdot]^{\prime})$ be the $2$-step nilpotent Lie algebra constructed as above. $\mathfrak{g}$ carries a natural almost complex structure $J_{\alpha}^{\prime}=J_{\alpha}\oplus I_{\alpha}$, where $\{I_{\alpha}\}$ is the standard hypercomplex structure on $\mathbb{R}^{4k}$. The following result is a characterization of $N_{J^{\prime}_{\alpha}}=0$.
@@
Let us now assume that the $2$-step nilpotent hypercomplex Lie algebra $(\mathfrak{n},\{J_{\alpha}\})$ satisfies that each $J_{\alpha}$ is $2$-step. If there exists $X\in\mathfrak{n}^{1}$ such that $\mu(J_{\alpha}X,\cdot)\neq 0$, then the corresponding complex structure $J^{\prime}_{\alpha}$ on $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k}$ is $3$-step, according to Lemma 2.4. We will use this procedure to construct new examples of $2$-step nilpotent hypercomplex Lie algebras where one, two or all three complex structures are $3$-step.
@@
Let $\mathfrak{n}_{8}$ be the $8$-dimensional 22-step nilpotent Lie algebra
$de^{i}=0,\ \ i=1,\dots,7,\ \ de^{8}=e^{12}-e^{34}.$ (4.2)
It has been shown in [DF1, DF2] that $\mathfrak{n}_{8}$ admits a hypercomplex structure $\{J_{\alpha}\}$ defined as
@@
$\mu=e^{15}\otimes e_{9}+e^{25}\otimes e_{10}+e^{35}\otimes e_{11}+e^{45}\otimes e_{12}.$
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{15},\ \ de^{10}=-e^{25},\\ &de^{11}=-e^{35},\ \ de^{12}=-e^{45},\end{split}$
@@
$\mu=\big{(}e^{15}+e^{16}\big{)}\otimes e_{9}+\big{(}e^{25}+e^{26}\big{)}\otimes e_{10}+\big{(}e^{35}+e^{36}\big{)}\otimes e_{11}+\big{(}e^{45}+e^{46}\big{)}\otimes e_{12}.$
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{15}-e^{16},\ \ de^{10}=-e^{25}-e^{26},\\ &de^{11}=-e^{35}-e^{36},\ \ de^{12}=-e^{45}-e^{46},\end{split}$
@@
$\mu=e^{56}\otimes e_{9}-e^{67}\otimes e_{10}+e^{57}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. Since $\mu(J_{1}e_{8},\cdot)=\mu(-e_{7},\cdot)\neq 0$$\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent complex structures.
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{12}-e^{34},\\ &de^{9}=-e^{56},\ \ de^{10}=0,\ \ \ de^{11}=e^{67},\ \ de^{12}=-e^{57},\end{split}$
@@
We point out that the $2$-step nilpotent Lie groups corresponding to the three Lie algebras constructed in this section admit lattices [Mal].
@@
We present next a construction which is a slight generalization of [BF, Lemma 3.1].
@@
The Jacobi identity on $\mathfrak{h}$ follows from the fact that $\rho$ is a Lie algebra homomorphism. Note that $\{(0,V):V\in\mathbb{R}^{4k}\}$ is an abelian ideal of $\mathfrak{h}$. Let $\widetilde{J}_{\alpha}$, $\alpha=1,2,3$, be the endomorphisms of $\mathfrak{h}$ defined by $\widetilde{J}_{\alpha}=J_{\alpha}\oplus I_{\alpha}$.

### v2

@@ Abstract
We study the holonomy of the Obata connection on 2-step hypercomplex nilmanifolds. By explicitly computing the curvature tensor, we determine the conditions under which the Obata connection is flat, showing that this depends on the nilpotency step of each complex structure. In particular, we show that for 2-step hypercomplex nilmanifolds the holonomy algebra of the Obata connection is always an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$ and we prove that the $\mathbb{H}$-solvable conjecture holds in this case. Furthermore, we provide new examples of $k$-step nilpotent hypercomplex nilmanifolds, with arbitrary $k$, which are not Obata flat.
@@ 1 Introduction
Nilmanifolds, that is, compact quotients $\G\Gamma\backslash$ G of connected and simply connected nilpotent Lie groups $G$ by discrete subgroups $\Gamma$, have played a fundamental role in differential geometry and topology since the seminal works of Mal’cev and Mostow. They provide a natural class of spaces where algebraic and geometric properties interact tightly, offering a fair balance between algebraic manageability and rich geometric behaviour. In low dimensions, the latter allows classification results for various geometric structures (e.g., [Sal, FPS, Ug]), especially when these are left-invariant.
@@
In the hypercomplex setting, many of these nice behaviours fail to carry over in a unified way. Each complex structure $J_{\alpha}$ ($\alpha=1,2,3$) may satisfy suitable compatibility conditions individually, but it is not clear whether these structures collectively induce analogous global properties. For instance, it is natural to ask whether the subspace
@@
A central geometric object in this setting is the Obata connection: the unique torsion-free connection that preserves all three complex structures. Its holonomy group $\operatorname{Hol}(\nabla)$ lies in $\operatorname{GL}(n,\mathbb{H})$ and plays a crucial role in understanding the geometry of hypercomplex manifolds [IP, SV]. In general, explicit descriptions of $\operatorname{Hol}(\nabla)$ are rare (see, for instance, [Sol, AGT, BFGV]). For nilmanifolds, however, it is known that the holonomy is contained in the commutator subgroup $\operatorname{SL}(n,\mathbb{H})$ [BDV, Corollary 3.3].
@@
This naturally leads to the question: Are all 2-step nilpotent hypercomplex Lie algebras Obata-flat? We answer this in the negative and clarify the role of the nilpotency step of each complex structure (see Sections 2 and 3 for details). In particular, we show that if the Obata connection is not flat, then each $J_{\alpha}$ must be 3-step nilpotent. No such examples had been previously known.
@@
Turning to structural properties, we show the Gorgynian conjecture for $2$-step hypercomplex Lie algebras (Theorem 5.4).
In Section 6, we prove that the holonomy algebra of the Obata connection is always abelian for $2$-step nilpotent hypercomplex Lie groups, and we describe it explicitly (Theorem 6.1). This reveals a surprising rigidity: even when the Obata connection is non-flat, the holonomy remains severely restricted. We further show that the Obata holonomy group of a hypercomplex nilmanifold which is at most 2-step is trivial if and only if the nilmanifold is a torus (Theorem 6.9).
In the final part of the paper, we extend our analysis to the $3$-step nilpotent case, showing that, under suitable conditions, the holonomy of the Obata connection remains abelian (Proposition 7.2).
Acknowledgments. A. Andrada and M. L. Barberis were partially supported by CONICET, SECyT-UNC and ANPCyT (Argentina). B. Brienza was partially supported by Project PRIN 2022 “Real and complex manifolds: Geometry and Holomorphic Dynamics”, by GNSAGA (Indam) and by the INdAM - GNSAGA Project CUP E53C24001950001. The third author is grateful to A. Fino, G. Gentili, G. Gorginyan, and M. Verbitsky for many helpful discussions, and to the FAMAF for the warm hospitality during the preparation of this work. The authors would like to thank the anonymous referee for their detailed review of the manuscript.
@@ 2 Preliminaries
It is known that $(M,\{J_{\alpha}\})$ is quaternionic (i.e., locally modelled over $\mathbb{H}^{n}$) if and only if the Obata connection $\nabla$ is flat [Ob, Theorem 11.2]. This highlights a fundamental difference with the complex case. For instance, the $K3$ surfaces and the homogeneous spaces $\operatorname{SU}(3)$ and $\operatorname{SO}(6)/\operatorname{SU}(2)$ provide examples of hypercomplex but not quaternionic manifolds [Boy, Joy].
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}$. A hypercomplex structure $\{J_{\alpha}\}$ on $G$ is said to be left-invariant if left translations by elements of $G$ are hyperholomorphic, i.e., holomorphic with respect to each $J_{\alpha}$. In this case, the structure is completely determined by its value at the identity, which corresponds to a hypercomplex structure on $\mathfrak{g}$.
If $\Gamma$ is a lattice in $G$, any left-invariant hypercomplex structure on $G$ descends to a hypercomplex structure on the compact quotient $\G\Gamma\backslash$ G, which is then called invariant. In this setting, the natural projection $\GG\to\Gamma\backslash G$ is hyperholomorphic.
@@ A Lie algebra $\mathfrak{g}$ is said to be nilpotent if its lower central series
Remark 2.2.
Equivalently, a Lie algebra $\mathfrak{g}$ is nilpotent if its upper central series
$0=\mathfrak{g}_{0}\subset\mathfrak{g}_{1}\subset\cdots\subset\mathfrak{g}_{k}\subset\cdots,$
defined recursively by
$\mathfrak{g}_{k}=\{X\in\mathfrak{g}\mid[X,\mathfrak{g}]\subset\mathfrak{g}_{k-1}\},\qquad k\geq 1,$
eventually coincides with $\mathfrak{g}$. Moreover, the smallest integer $k$ such that $\mathfrak{g}_{k}=\mathfrak{g}$ coincides with the nilpotency step defined via the lower central series.
@@
It is well known that the existence of a lattice for a connected and simply connected nilpotent Lie group $N$ is equivalent to the existence of a rational basis of its Lie algebra $\mathfrak{g}$ [Mal], that is, a basis of $\mathfrak{g}$ for which the structure constants are rational numbers.
@@
The ascending series $\{\mathfrak{a}_{k}\}_{k\geq 0}$ (compatible with $J$) of $\mathfrak{g}$ is defined inductively by
$\mathfrak{a}_{0}=0,\quad\mathfrak{a}_{k}:=\{X\in\mathfrak{g}\ |\ [X,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\ \ \text{and}\ \ [JX,\mathfrak{g}]\subset\mathfrak{a}_{k-1}\},\ k\geq 1.$ (2.4)
@@ 3. $(3)$
Remark 2.6.
Recall that a complex structure $J$ on a Lie algebra $\mathfrak{g}$ is called abelian if it satisfies $[JX,JY]=[X,Y]$ for all $X,Y\in\mathfrak{g}$ (see [BDM]). In the case when $\mathfrak{g}$ is 2-step nilpotent, an abelian complex structure is $2$-step nilpotent, since the center is $J$-invariant.
A hypercomplex structure $\{J_{\alpha}\}$ on a Lie algebra $\mathfrak{g}$ is called abelian if $J_{\alpha}$ is abelian for some $\alpha=1,2,3$, and it follows from [DF2, Lemma 3.1] that $J_{y}$ is abelian for each $y\in S^{2}$.
@@ 3 Computation of the Obata curvature
where the second equality follows from the integrability of $J_{\alpha}$ on the first term and the integrability of $J_{\gamma}$ on the second term. ∎ As a consequence of Proposition 3.1 we obtain the following result. This is a generalization of [B1, Proposition 5.1], where it was shown that if $\mathfrak{z}$ is $J_{\alpha}$ invariant for all $\alpha$ then $\{J_{\alpha}\}$ is Obata flat.
@@
If $J_{\beta}$ is $2$-step nilpotent then $J_{\beta}\mathfrak{g}^{1}\subset\mathfrak{z}$ (Lemma 2.5). The result then follows from Proposition 3.1, since each term contains an element of the form $\cdot,J_{\beta}\mathfrak{g}^{1}]$. ∎
@@ 4 Construction of non-flat 2-step hypercomplex nilmanifolds
To the best of our knowledge, in all known examples of $2$-step nilpotent hypercomplex Lie algebras $(\mathfrak{g},\{J_{\alpha}\})$, each complex structure $J_{\alpha}$ is $2$-step [B2, DF1, DF3, BD, LW, FG]. In particular, every known example of a $2$-step nilpotent Lie group equipped with a left-invariant hypercomplex structure is Obata flat (see Corollary 3.2).
According to [DF3, Section 3], in dimension $8$ any nilpotent Lie algebra admitting a hypercomplex structure is $2$-step nilpotent, and, moreover, each complex structure $J_{\alpha}$ is $2$-step.
In this section, we construct examples of $2$-step nilpotent hypercomplex Lie algebras for which one, two, or all of the complex structures are $3$-step. Moreover, we exhibit an example in which all three complex structures are $3$-step and the associated Obata connection is not flat.
@@
Let $\mathfrak{n}$ be a $2$-step nilpotent Lie algebra and $\mathfrak{g}:=\mathfrak{n}\oplus\mathbb{R}^{r}$for some $r\in\mathbb{N}$.
Let $\mu\in\bigwedge^{2}\mathfrak{n}^{*}\otimes\mathbb{R}^{r}$ be a $2$-form on $\mathfrak{n}$ with values in $\mathbb{R}^{r}$ such that $\mathfrak{n}^{1}\subset\ker(\mu)$, where $\mathfrak{n}^{1}=[\mathfrak{n},\mathfrak{n}]$. We use $\mu$ to define a bracket $\cdot$$\cdot$$^{\prime}$ on $\mathfrak{g}$ as follows:
@@
Then, $(\mathfrak{g},[\cdot,\cdot]^{\prime})$ is a $2$-step nilpotent Lie algebra. We note that the condition $\mathfrak{n}^{1}\subset\ker(\mu)$ guarantees that the bracket $\cdot$$\cdot$$^{\prime}$ satisfies the Jacobi identity and that $\mathfrak{g}$ is $2$-step nilpotent.
Let us now assume that $\mathfrak{n}$ is a $2$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$, and let $(\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k},[\cdot,\cdot]^{\prime})$ be the $2$-step nilpotent Lie algebra constructed as above. Then $\mathfrak{g}$ carries a natural almost complex structure $J_{\alpha}^{\prime}=J_{\alpha}\oplus I_{\alpha}$, where $\{I_{\alpha}\}$ is the standard hypercomplex structure on $\mathbb{R}^{4k}$. The following result is a characterization of $N_{J^{\prime}_{\alpha}}=0$.
@@
Let us now assume that the $2$-step nilpotent hypercomplex Lie algebra $(\mathfrak{n},\{J_{\alpha}\})$ satisfies the condition that each $J_{\alpha}$ is $2$-step. If there exists $X\in\mathfrak{n}^{1}$ such that $\mu(J_{\alpha}X,\cdot)\neq 0$, then the corresponding complex structure $J^{\prime}_{\alpha}$ on $\mathfrak{g}=\mathfrak{n}\oplus\mathbb{R}^{4k}$ is $3$-step (Lemma 2.5). Since $X\in\mathfrak{n}^{1}$ we may write
$X=\sum_{j}\,[Y_{j},Z_{j}],\ \ \ Y_{j},Z_{j}\in\mathfrak{n}.$
We define
$X^{\prime}:=\sum_{j}\,[Y_{j},Z_{j}]^{\prime}=\sum_{j}([Y_{j},Z_{j}],\mu(Y_{j},Z_{j}))\in\mathfrak{g}^{1}.$
We claim that $J_{\alpha}^{\prime}X^{\prime}\notin\mathfrak{z}(\mathfrak{g})$. In fact, let $W\in\mathfrak{n}$ be such that $\mu(J_{\alpha}X,W)\neq 0$, then
$J^{\prime}_{\alpha}X^{\prime},W]^{\prime}=(0,\mu(J_{\alpha}X,W))\neq 0.$
We will use this procedure to construct new examples of $2$-step nilpotent hypercomplex Lie algebras where one, two or all three complex structures are $3$-step.
@@
All $8$-dimensional hypercomplex nilpotent Lie algebras were classified in [DF3]. We consider the $8$-dimensional 22-step nilpotent Lie algebra $\mathfrak{n}_{8}$ defined by  [^1] We adopt the following notation: $e^{i,j}=e^{i}\wedge e^{j}$.
$de^{i}=0,\quad i=1,\dots,7,\qquad de^{8}=e^{1,2}-e^{3,4},$ (4.2)
which corresponds to the Lie algebra $\mathfrak{n}_{1}$ in [DF3, Section 3]. It was shown in [DF3, Section 3] (see also [DF1, Section 5]) that $\mathfrak{n}_{8}$ admits a hypercomplex structure $\{J_{\alpha}\}$, defined as follows:
@@
$\mu=e^{1,5}\otimes e_{9}+e^{2,5}\otimes e_{10}+e^{3,5}\otimes e_{11}+e^{4,5}\otimes e_{12}.$
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{1,5},\ \ de^{10}=-e^{2,5},\\ &de^{11}=-e^{3,5},\ \ de^{12}=-e^{4,5},\end{split}$
@@
$\mu=\big(e^{1,5}+e^{1,6}\big)\otimes e_{9}+\big(e^{2,5}+e^{2,6}\big)\otimes e_{10}+\big(e^{3,5}+e^{3,6}\big)\otimes e_{11}+\big(e^{4,5}+e^{4,6}\big)\otimes e_{12}.$
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{1,5}-e^{1,6},\ \ de^{10}=-e^{2,5}-e^{2,6},\\ &de^{11}=-e^{3,5}-e^{3,6},\ \ de^{12}=-e^{4,5}-e^{4,6},\end{split}$
@@
$\mu=e^{5,6}\otimes e_{9}-e^{6,7}\otimes e_{11}+e^{5,7}\otimes e_{12}.$
Since $\mu(e_{8},\cdot)=0$, and Equation 4.1 is satisfied for any $\alpha=1,2,3$, the $2$-step nilpotent Lie algebra $(\mathfrak{g}=\mathfrak{n}_{8}\oplus\mathbb{R}^{4},[\cdot,\cdot]^{\prime})$ can be endowed with the hypercomplex structure $\{J^{\prime}_{\alpha}=J_{\alpha}\oplus I_{\alpha}\}$. Since $\mu(J_{1}e_{8},\cdot)=\mu(-e_{7},\cdot)\neq 0$, $\mu(J_{2}e_{8},\cdot)=\mu(e_{6},\cdot)\neq 0$ and $\mu(J_{3}e_{8},\cdot)=\mu(-e_{5},\cdot)\neq 0$, $J^{\prime}_{1},J^{\prime}_{2}$ and $J^{\prime}_{3}$ are $3$-step nilpotent complex structures.
@@
$\begin{split}&de^{i}=0,\ \ \ i=1,\dots,7,\ \ \ de^{8}=e^{1,2}-e^{3,4},\\ &de^{9}=-e^{5,6},\ \ de^{10}=0,\ \ \ de^{11}=e^{6,7},\ \ de^{12}=-e^{5,7},\end{split}$
@@
We point out that the $2$-step nilpotent Lie groups corresponding to the three Lie algebras constructed in this section admit lattices, since all the bases considered in the examples are rational.
@@
Next, we present a construction which is a slight generalization of [BF, Lemma 3.1].
@@
The Jacobi identity on $\mathfrak{h}$ follows from the fact that $\rho$ is a Lie algebra homomorphism. Note that $\{(0,V)\mid V\in\mathbb{R}^{4k}\}$ is an abelian ideal of $\mathfrak{h}$. Let $\widetilde{J}_{\alpha}$, $\alpha=1,2,3$, be the endomorphisms of $\mathfrak{h}$ defined by $\widetilde{J}_{\alpha}=J_{\alpha}\oplus I_{\alpha}$.